A scuba divers oxygen tank reads 900 mmHg while on her boat, where the temperature is 27 degrees C. When she dives down to the bottom of the lake and experiences a change is atmospheric pressure the temperature will drop down to –183 degrees C. What will the pressure in her backup tank be at that temperature? (270 mmHg)
1
Expert's answer
2020-05-16T13:44:16-0400

We know that for constant volume


"\frac{P_2}{T_2}=\frac{P_1}{T_1}"

"P_1= 900\ mm\ Hg\\T_1=300\ K\\T_2=90\ K"

So

"P_2=\frac{P_1T_2}{T_1}=\frac{900\times90}{300}= 270\ mm\ Hg"